Windsor, Inc. purchased a 777 aircraft on January 1, 2020 at a cost of $40,280,000. The estimated useful life of the aircraft is 20 years, with an estimated salvage value of $6,042,000.

Compute the accumulated depreciation and book value at December 31, 2022 using the straight-line method and the double-declining-balance method.

Straight-line Declining-balance

Accumulated depreciation

$enter a dollar amount $enter a dollar amount

Book value

$enter a dollar amount $enter a dollar amount
Answer #1

Straight-line method

Annual depreciation expense = (Cost of equipment - Residual value)/Useful life

= (40,280,000 - 6,042,000)/20

= 34,238,000/20

= $1,711,900

Accumulated depreciation for 3 years = Annual depreciation expense x 3

= 1,711,900 x 3

= $5,135,700

Book value of equipment at December 31, 2022 = Cost of equipment - Accumulated depreciation for 3 years

= 40,280,000 - 5,135,700

= $35,144,300

Double-declining-balance method

Double declining depreciation rate = 2 x 1/Useful life

= 2 x 1/20

= 10%

Depreciation schedule

Year

Beginning value

Depreciation expense

Accumulated depreciation

Ending value

2020

40,280,000

40,280,000 x 10% = 4,028,000

4,028,000

40,280,000 - 4,028,000 = 36,252,000

2021

36,252,000

36,252,000 x 10% = 3,625,200

4,028,000 + 3,625,200 = 7,653,200

36,252,000 - 3,625,200 = 32,626,800

2022

32,626,800

32,626,800 x 10% = 3,262,680

7,653,200 + 3,262,680 = $10,915,880

3,262,680 - 3,262,680 = 29,364,120

Straight-line Declining-balance

Accumulated depreciation

$5,135,700 $10,915,880

Book value

$35,144,300 $29,364,120